OpenStudy (anonymous):

a cone with slant height 22 cm and radius 7 cm about 484 cm2 about 507 cm2 about 637 cm2 about 1,123 cm2

OpenStudy (anonymous):

If you need the surface area, remember that SA = pi*(r^2) + pi*(r)*(h) = pi*(r^2 + r*h) = pi*(49 + 154) = 203*pi that's closer to 637cm^2 than all alternatives. If you don't need the SA, I misinterpreted your question, I am sorry.

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

You are welcome :-)